# A multi-band study of pulsar glitches with Fermi-LAT and Parkes

### v1

@@ 4.2 PSR J1028--5819
According to existing studies, large glitches with exponential recovery are often accompanied by long-term linear recovery, which continues until the next glitch event occurs (Yu et al., 2013). The large glitch in PSR J1028--5819 follows this trend. As shown in panel (c) of Fig. 2, the blue line represents the fitted result for the linear decrease in $\dot{\nu}$. We found that the post-glitch slope ($\ddot{\nu}\sim 2.14(5)\times 10^{-23}$ s-3) is nearly three times larger than the pre-glitch slope ($\ddot{\nu}\sim 7.35(1)\times 10^{-24}$ s-3), indicating that the glitch caused a permanent change in the spin-down rate.
We fit this glitch with Eqs. (7) and (14), and the fit result is shown in Fig. 2. The total fractional crustal superfluid moment of inertia involved in the glitch is estimated to be $I_{\rm cs}/I=3.35(16)\%$, as seen in Table 5. The vortex creep model predicts that the time to the next glitch following G1 is $t_{\rm th}=4890(269)$ d. Since not all of the angular momentum reservoir within the crustal superfluid was depleted during G1, PSR J1028--5819 may experience another glitch due to a different pinned superfluid layer before the recovery from the previous event is fully complete. The fit result for the crustal superfluid’s recoupling timescale, $\tau_{\rm nl}=71$ d, is close to the lower end of the theoretical estimate range $\tau_{\rm nl}=$ 108–5113 d, as derived from Eq. (15).
@@
Figure 4: PSR J1709--4429 timing residuals of glitch 4. The three panels represent the timing residuals for the spin-down model with 0, 1, and 2 exponential decay terms, respectively. The vertical gray dashed line and the number of parentheses at the top represent the glitch epoch and the reference number of glitch occurrences, respectively.
@@ 4.3 PSR J1709--4429 (B1706--44)
In 1992, the pulsed radio emission from PSR J1709--4429 was discovered for the first time in the Parkes high-frequency survey project (Johnston et al., 1992). Thompson et al. (1996) detected gamma-ray pulsations with EGRET, and Chandra found X-ray pulsed radiation in 2002 (Gotthelf et al., 2002). This is a bright Vela-like pulsar with a period of 102.5 ms, characteristic age of 17.5 kyr and the spin-down energy loss rate of $\dot{E}\sim 3.4\times 10^{36}$ erg s-1 (Wang et al., 2000). In 1995, the glitch event of PSR J1709--4429 was reported for the first time (Johnston et al., 1995), and five glitches were successively published in the following timing observations (Weltevrede et al., 2010; Yu et al., 2013; Lower et al., 2018, 2021).
@@
For glitch 4, Yu et al. (2013) reported that post-glitch behaviour has only one exponential recovery. However, when we perform a fit combining Fermi-LAT and Parkes timing data, two exponential decay characteristics are revealed in the timing residuals, as shown in Fig. 4. The first exponential decay feature is $Q$1 = 0.0104(5) and $\tau_{\rm d1}$ = 72(4) d, and the second exponential decay feature is $Q$2 = 0.006(1) and $\tau_{\rm d2}$ = 4.2(6) d. Thus, the transient increment in the frequency and frequency first derivative is $\Delta\nu/\nu\sim 2777(4)\times 10^{-9}$ and $\Delta\dot{\nu}/\dot{\nu}\sim 64(15)\times 10^{-3}$, respectively.
For glitches 5 and 6, the glitch parameters obtained by our fitting are consistent with the previously reported results within the error range, and the detailed results are shown in Table 3. We performed exponential and linear fits to the evolution of $\dot{\nu}$, with the results displayed as red and blue dashed lines, respectively, in panel (c). For glitch 4, both single and double exponential functions were used to model the post-glitch exponential recovery of $\dot{\nu}$. The accuracy of fit was evaluated using the coefficient of determination ($R^{2}=1-\frac{\sum_{\rm i}(Y_{\rm i}-\overline{y})^{2}}{\sum_{\rm i}(y_{\rm i}-\overline{y})^{2}}$), where $y_{\rm i}$ represents the data points, $\overline{y}$ is their mean, and $Y_{\rm i}$ is the fitted value.
By fitting the exponential recovery of $\dot{\nu}$, we found that glitch 4 achieved the highest accuracy of fit ($R^{2}=0.95$) with a double exponential model, compared to $R^{2}=0.70$ for a single exponential fit. The result is consistent with that of the two exponential recovery procedures observed in the data processing. For glitches 5 and 6, an exponential term accurately describes the recovery of $\dot{\nu}$, with the fit achieving an accuracy of over 0.99. Interestingly, although the linear recovery slopes change significantly after glitches 1, 2, and 3 (Yu et al., 2013), our linear fitting results show that the linear recovery process after glitches 4, 5, and 6 exhibits a similar slope, which is approximately $\ddot{\nu}\sim 2.0\times 10^{-22}$ s-3.
We apply the vortex creep model to fit the three glitches, using Eqs. (7) and (14), and the fit result is shown in Fig. 3 panel (d):
  • G4: The exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=86(9)$ d is slightly different than $\tau_{\rm d}=72(4)$ d found from fit with Eq. (2) (see Table 3) again due to the initial exponential dependency of Eq. (14).  [^7] We mention that, alternatively, the $\tau_{\rm d}=4.2$ d time-scale exponential recovery associated with G4 of PSR J1709--4429 either arises from shallow densities of the crust (according to the vortex resistive model) or intermediate densities of the core region (according to the vortex cluster model). See more related discussions in Sec. 5.3. The vortex creep model prediction for the time to the next glitch from G4 is $t_{\rm th}=1680(16)$ d and in qualitative agreement with the observed inter-glitch time $t_{\rm obs}=1648(4)$ d. The total fractional crustal superfluid moment of inertia involved in G4 after corrected for the crustal entrainment effect is $I_{\rm cs}/I=2.30(2)\%$. The recoupling time-scale for the crustal superfluid associated with G4 is $\tau_{\rm nl}=3.4$ d.
@@ 4.4 PSR J1718--3825
PSR J1718--3825 is a young pulsar, first discovered by Manchester et al. (2001) in the radio band. It has a characteristic age of 89.5 kyr, a spin period of 74.7 ms, and a relatively high spin-down luminosity of $\dot{E}\sim 1.3\times 10^{36}$ erg s-1. In addition to an associated X-ray nebula (Hinton et al., 2007), it is also a source of gamma-ray emission (HESS J1718) within the supernova remnant RX J1713.7--3946 (Aharonian et al., 2007).
We analysed more than 14 yr of timing data for PSR J1718--3825 in the radio and gamma-ray bands. A total of 3 glitches were identified, of which glitch 1 and 2 have been reported in the radio band (Yu et al., 2013; Lower et al., 2021), and glitch 3 is a new small glitch. By examining changes in the timing residuals, it was deduced that glitch 3 occurred at MJD $\sim$ 59121(8). Using the fitting method, we determine the fractional glitch size to be $\Delta\nu/\nu\sim 1.9(2)\times 10^{-9}$ and the first frequency derivative to be $\Delta\dot{\nu}/\dot{\nu}\sim-0.12(9)\times 10^{-3}$. We also used the pre- and post-glitch timing data of glitch 2 to analyse the timing noise, as the time span exceeds three years. It was found that the spectral index before and after glitch 2 remained consistent at 6.
In Fig. 5, the three detected glitches in PSR J1718--3825 are presented. It is clear that $\dot{\nu}$ only experienced a long-term linear recovery after the glitches, and there may be stable $\ddot{\nu}$. Therefore, based on the glitch epochs, the $\dot{\nu}$ data were divided into four intervals and linear fits were applied to each interval to determine $\ddot{\nu}$. The fits revealed that, under the influence of glitch, $\ddot{\nu}$ has undergone obvious changes, with values of $2.0(3)\times 10^{-23}$ s-3, $2.887(3)\times 10^{-23}$ s-3, $1.65(3)\times 10^{-23}$ s-3, and $2.61(4)\times 10^{-23}$ s-3, respectively. These fitted values are represented by the blue dashed lines in panel (c) of Fig. 5. A comparison with predictions from the superfluid vortex bending model is also provided for the first (G1) and second (G2) glitch events. Detailed discussions are included below in Sec. 5.
@@ 4.5 PSR J1509--5850
PSR J1509--5850 is a middle-aged pulsar with a characteristic age of 154 kyr (Kramer et al., 2003). We have analysed nearly 14 yr of timing data in both the radio and gamma-ray bands and found no glitches within the observed data range. Our study of the spin-down rate evolution suggests that PSR J1509--5850 exhibits a roughly stable $\ddot{\nu}$, as shown in Fig. 6. Using the least squares method for linear fitting, we determined $\ddot{\nu}$ to be $1.328(4)\times 10^{-24}$ s-3, with the fitted result indicated by the blue dotted line in Fig. 6. Additionally, the figure presents a fit based on the superfluid vortex bending model, which describes well the $\dot{\nu}$’s oscillation (see Sec. 5).
@@ 5.1 Oscillations in the spin-down rate
We apply the vortex bending model to the timing solution of PSR J1509--5850 between MJD 54303–59510, see Fig. 6. The model free parameters corresponding to the vortex bending model’s application are given in Table 6. If we attribute the large frequency second time derivative of $\ddot{\nu}\cong 1.37\times 10^{-24}$  s-3 to an unresolved Vela-like glitch occurred before the time interval we considered, then Eq. (24) with the typical value $\Delta\dot{\nu}_{\rm p}/\dot{\nu}=5\times 10^{-3}$ gives the estimate of $t_{\rm ig}=158$ yr for the time between two successive giant glitches hypothesised to be experienced by this pulsar. Thus, we predict that PSR J1509--5850 will glitch again within the same timescale. Future observations of this pulsar are required to test our assertion.
We also analyse the timing data of PSR J1718--3825 between MJD 54303–59491. Since there is not enough data for the third glitch (G3), we made an application of the vortex bending model to its first (G1) and second (G2) glitch events. See Fig. 5, panel (c). The fit parameters corresponding to the vortex bending model are given in Table 6. While the vortex bending model slightly improves the fit to the data of G1 compared to the linear one, it provides an excellent fit to the G2 data reducing the residuals significantly over the linear one. In particular, the damping timescale of the oscillations in the spin-down rate is well constrained to be $\tau=218$ d. When assessed with Eq. (8), this would provide important aspects for the microphysical parameters related to nuclear pinning mechanism as well as temperature estimate for the neutron star crust. Also, the length-scale $\ell$ found from the fits would provide a new method to put constraints on the crustal thickness and in turn on the EOS of the neutron stars. It should be also important to address these constraints in light of the sinusoidal-like oscillations observed in the spin-down rates of various pulsars.
@@ 5.2 Inter-glitch Braking Index
If only the external torque is considered, based on the use of torque laws of the form $-K\Omega^{n}$, where $K$ is a constant and the braking index $n={\nu\ddot{\nu}}/{\dot{\nu}}^{2}$ is assumed to be 3 as predicted by the magnetic dipole model (Manchester & Taylor, 1977). Following the vortex creep model (Alpar et al., 1984; Gügercinoǧlu & Alpar, 2017), glitches may occur as a result of collective unpinning of a very large number of vortex lines in the crustal superfluid and inter-glitch spin evolution of a pulsar is dominated by the response of the star to the variations in the superfluid internal torques’ coupling due to such vortex discharges. In terms of the model parameters theoretical expectation of the inter-glitch braking index can be recast as (Alpar & Baykal, 2006):
@@
where the glitch magnitudes in the rotation and the spin-down rates are expressed in units of $10^{-6}$ and $10^{-3}$, respectively. Also, we have used the shorthand notation $\beta=I_{\rm B}/I_{\rm A}$. As a fiducial value, in our calculations below we take $\beta=4.11$, which was obtained for PSR J1048–5832 (Liu et al., 2024). In comparison, the $constant$ braking index ($n$) before and after the glitch were estimated according to the fitted $\ddot{\nu}$ and the parameters ${\nu}$, $\dot{\nu}$ in Table 2. We observed stable values of $\ddot{\nu}$ for PSRs J1028--5819, J1509--5850 and J1718--3825 and fitted the slopes of $\ddot{\nu}$ for each. In the case of PSR J1028--5819, the braking indices before and after the glitch are 21.66(4) and 62.8(4), respectively. For glitch 2 of PSR J1718--3825, the braking indices before and after the glitch are 69.35(7) and 41.0(8), respectively. Although we did not identify a glitch event for PSR J1509--5850, its $\dot{\nu}$ has steadily evolved over 14 yr, resulting in a braking index of $n=11.12(2)$. These non-canonical braking indices may be indicative of internal torque at work, likely associated with the dynamics of nucleon superfluidity in this context.
Moreover, the most prominent effect of vortex unpinning on the spin parameters is $\ddot{\nu}$ and given by
@@
$t_{\rm ig}=\frac{\Delta\dot{\nu}_{\rm p}}{\ddot{\nu}_{\rm ig}}=2\times 10^{-3}\left(\beta+\frac{1}{2}\right)\tau_{\rm sd}\left(\frac{\Delta\dot{\nu}_{\rm p}}{\dot{\nu}}\right)^{-1}_{-3}\left(\frac{\Delta\nu_{\rm p}}{\nu}\right)_{-6}\ .$ (24)
@@
Comparison of the prediction of the vortex creep model [Eq. (23)] with observed inter-glitch $\ddot{\nu}$ for our sample of glitches is given in Table 7. An inspection to the Table 7 reveals that the agreement between the vortex creep model predictions and observations are qualitatively good except for G8 of PDR J1420--6048. The reason for the discrepancy may be a persistent step increase that not resolved which leads to a somewhat larger estimate in $\Delta\dot{\nu}_{\rm p}$. Future observations will help to test this hypothesis.
5.3 Exponential Recovery Mechanisms
Figure 7: Exponential decay time-scale of the the neutron star core predicted by the vortex cluster model [Eq. (25)] (in logarithmic scale) versus matter density for PSR J1028--5819 (blue dotted curve), PSR J1420--6048 (green solid curve) and PSR J1709--4429 (red dashed curve).
For glitch events with large frequency changes, exponential recovery may occur after glitch (Yu et al., 2013; Lower et al., 2021; Basu et al., 2022). The exponential recovery timescale is generally between 10 and 300 d, with a longer timescale for older pulsars (Yu et al., 2013). There may be more than one exponential recovery process after the glitches, for example, Vela pulsar has been detected four exponential recovery processes in its glitch 14 (Yu et al., 2013; Dodson et al., 2002). In the vortex creep model discussed above, the outer core superfluid’s response to a glitch involves exponential relaxation through vortex creep across magnetic flux tubes. The corresponding decay timescale has been calculated for several gamma-ray glitching pulsars (Gügercinoğlu et al., 2022; Liu et al., 2024). Considering the complex dynamical behaviour observed in the data it would be worthwhile to explore alternatives that differ from the vortex creep model in several important details. Here, we consider two alternative mechanisms that could explain the post-glitch exponential recovery and apply them to the observational data for the sample analysed in this study.
Sedrakian & Sedrakian (1995) considered a different mechanism based on vortex line-flux tube interaction originated from the whole neutron star core. According to them, initial magnetic field of a neutron star inherited from its progenitor star is expelled out from the core region as a result of setting up of Meissner currents after transition to superconducting state. However, up to a certain distance from its centre known as the London penetration depth, each vortex line is attached with magnetic flux due to proton supercurrent circulating around them. Immediate vicinity of a vortex line is surrounded by more numerous flux tubes and a so-called vortex cluster is formed. Electrons’ scattering off such vortex clusters couples the superfluid/superconducting core of a neutron star to its crust and leads to exponential decay following glitches. Vortex cluster model predicts for the decay time-scale (Sedrakian & Sedrakian, 1995),
$\tau_{\rm d,cluster}=\frac{\sqrt{3}\pi}{2048}\frac{\hbar}{m_{\rm e}}\frac{\rho_{\rm e}}{\rho_{\rm s}^{*}}\frac{1}{k_{\rm e}\lambda_{\rm p}^{3}}\mu|k|\left(\frac{\xi_{\rm p}}{\lambda_{\rm p}}\right)^{\frac{1}{|k|}}P^{2}\ ,$ (25)
where $\hbar$ is the Planck constant divided by $2\pi$, $m_{\rm e}$ is the mass of the electron, $\rho_{\rm e}$ is the mass density of electrons, $\rho_{\rm s}^{*}$ is the mass density of the entrained superfluid neutrons, $k_{\rm e}$ is the electron wave number, $\lambda_{\rm p}$ is the London penetration depth for proton supercurrents, $\xi_{\rm p}$ is the coherence length for superconducting protons, $k\equiv(m_{\rm p}^{*}-m_{\rm p})/m_{\rm p}$ is the entrainment coefficient for the superconducting protons, $\mu$ is the magnetic permeability of vortex cluster and $P$ is the spin period of the pulsar. For calculating these parameters Sedrakian & Sedrakian (1995) employed moderately stiff AV14+UV EOS of Wiringa et al. (1988), and took microphysical quantities associated with superfluidity/superconductivity from Baldo et al. (1992). We adopted these values as given in tables 1 and 2 of their paper. In Fig. 7 we give the exponential decay timescale estimate Eq. (25) for the three glitching gamma-ray pulsars studied in this paper. Due to large variation of the model free parameters throughout of the core region, estimate Eq. (25) ranges from several minutes to decades for a given pulsar and in practice able to explain all exponential recovery time-scales seen across all glitching pulsar population. And, as the only parameter differing among different pulsars in their model is the pulsar spin period, which occupies a narrow parameter space, the curves in Fig. 7 are quite similar.
Figure 8: Exponential decay time-scale of the neutron star crust predicted by the vortex resistive model [Eq. (26)] (in logarithmic scale) versus matter density for PSR J1028--5819 (blue dotted curve), PSR J1420--6048 (green solid curve) and PSR J1709--4429 (red dashed curve).
Jones (1990) considered the energy transfer process associated with the superfluid vortex line past the crustal lattice and mutual friction coupling to the phonons. At relatively low velocity lag between the vortex and the lattice the interaction is mediated by excitation and propagation of sound waves. Jones (1993) argued that since the neutron star crustal lattice is described by a polycrystalline structure and vortex lines have different orientations with respect to the nuclear lattice, the pinning forces arising from the randomly distributed point defects of lattice around the axis of vortex may largely cancel each other, at least for a portion of the inner crust superfluid. So, the dominant force acting on a vortex moving through the inner crust is the resistive force due to energy carried over to the phonons by above mentioned mechanism. Such coupling leads to exponential decay when vortices unpinned and displaced with respect to the ionic crystal. Decay timescale is given by (Jones, 1990)
$\tau_{\rm Jones90}=\frac{1}{2\mathcal{B}\Omega}=\frac{16\pi^{3/2}}{3}\frac{\xi^{3}_{\rm n}}{a}\frac{Mc^{3}_{\rm s}}{E^{2}_{\rm p}}\frac{\kappa\rho_{\rm s}}{\Omega},$ (26)
where, $a$ is the lattice constant, $\xi_{\rm n}$ is the vortex core radius or neutron superfluid’s coherence length, $M$ is the total mass in atomic nuclei, $E_{\rm p}$ is the pinning energy between a nucleus and vortex, $\kappa$ is the vorticity, $\rho_{\rm s}$ is the superfluid neutron’s mass density and $\Omega=2\pi/P$ is the angular velocity of the neutron star. In the above expression, $c_{\rm s}$ is the lattice phonon velocity and defined as the ratio of the plasma frequency to the Debye wave number. On the other hand, $\mathcal{B}$ is the mutual friction parameter and allows for comparison of the strength of this mechanism with other processes prevailing inside neutron stars (Haskell et al., 2012). Microphysical parameters in Eq. (26) were calculated for five distinct densities in the crust considered in Negele & Vautherin (1973) and were given in table 1 of Jones (1990). In Fig. 8 we give the exponential decay timescale estimate Eq. (26) for the three glitching gamma-ray pulsars studied in this paper. Although Eq. (26) does not exhibit diversity as much as that predicted by Eq. (25) for relevant neutron star parameters, the exponential decay time for this process also varies greatly by three orders of magnitude. Again, the curves in Fig. 8 are very similar to each other because the spin period, which is the only variable quantity other than nuclear physics parameters, does not change much among pulsars considered in this study. Note that, the sharp changes of lines in the figure do not correspond to physical discontinuities, but are due to a low number of data points (only 5 for each curve).

### v2

@@ 4.2 PSR J1028--5819
Large glitches with exponential recovery are often accompanied by long-term linear recovery, which continues until the next glitch event occurs (Yu et al., 2013). The large glitch in PSR J1028--5819 follows this trend. As shown in panel (c) of Fig. 2, the blue and red dashed lines represent the fitting curve of $\dot{\nu}$ linear and exponential decay, respectively. Among them, the $\chi^{2}_{\rm r}$ values of linear fitting in pre- and post-glitch remain consistent with 1.06, corresponding to d.o.f. of 18 and 8, respectively. Furthermore, the $\chi^{2}_{\rm r}$ value for the exponential fitting of post-gltch is 1.60, corresponding to a d.o.f. of 5. We found that the post-glitch slope ($\ddot{\nu}\sim 2.14(5)\times 10^{-23}$ s-3) is nearly three times larger than the pre-glitch slope ($\ddot{\nu}\sim 7.35(14)\times 10^{-24}$ s-3), indicating that the glitch caused a permanent change in the spin-down rate.
We fit this glitch with Eqs. (7) and (14), and the fit result is shown in Fig. 2. The $\chi^{2}_{\rm r}$ value of this fitting is 0.22, corresponding to a d.o.f. of 13. The total fractional crustal superfluid moment of inertia involved in the glitch is estimated to be $I_{\rm cs}/I=3.35(16)\%$, as seen in Table 5. The vortex creep model predicts that the time to the next glitch following G1 is $t_{\rm th}=4890(269)$ d. Since not all of the angular momentum reservoir within the crustal superfluid was depleted during G1, PSR J1028--5819 may experience another glitch due to a different pinned superfluid layer before the recovery from the previous event is fully complete. The fit result for the crustal superfluid’s recoupling timescale, $\tau_{\rm nl}=71$ d, is close to the lower end of the theoretical estimate range $\tau_{\rm nl}=$ 108–5113 d, as derived from Eq. (15).
@@
Figure 4: PSR J1709--4429 timing residuals of glitch 4. The three panels represent the timing residuals for the spin-down model with 0, 1, and 2 exponential decay terms, respectively. The vertical gray dashed line and the numbers in between parentheses at the top represent the glitch epoch and the reference number of glitch occurrences, respectively.
@@ 4.3 PSR J1709--4429 (B1706--44)
In 1992, the pulsed radio emission from PSR J1709--4429 was discovered in the Parkes high-frequency survey project (Johnston et al., 1992). Thompson et al. (1996) detected gamma-ray pulsations with EGRET, and Chandra found X-ray pulsed radiation in 2002 (Gotthelf et al., 2002). This is a bright Vela-like pulsar with a period of 102.5 ms, characteristic age of 17.5 kyr and the spin-down energy loss rate of $\dot{E}\sim 3.4\times 10^{36}$ erg s-1 (Wang et al., 2000). In 1995, the first glitch of PSR J1709--4429 was reported (Johnston et al., 1995), and five more glitches were successively published in the following timing observations (Weltevrede et al., 2010; Yu et al., 2013; Lower et al., 2018, 2021).
@@
Moreover, we performed exponential and linear fits to the evolution of $\dot{\nu}$, with the results displayed as red and blue dashed lines, respectively, in panel (c). These fitted $\chi_{\rm r}^{2}$ values range from 0.32 to 1.16, corresponding to a d.o.f. range of 4 to 12. For glitch 4, both single and double exponential functions were used to model the post-glitch exponential recovery of $\dot{\nu}$. By fitting the exponential recovery of $\dot{\nu}$, we found that glitch 4 achieved the highest accuracy of fit ($\chi_{\rm r}^{2}=0.32$, for 4 d.o.f.) with a double exponential model, compared to $\chi_{\rm r}^{2}=1.37$ (7 d.o.f.) for a single exponential fit. The result is consistent with that of the two exponential recovery procedures observed in the data processing. For glitches 5 and 6, an exponential term accurately describes the recovery of $\dot{\nu}$, with $\chi_{\rm r}^{2}$ values close to 1, specifically 0.81 (7 d.o.f.) and 1.12 (5 d.o.f.). Interestingly, although the linear recovery slopes change significantly after glitches 1, 2, and 3 (Yu et al., 2013), our linear fitting results show that the linear recovery process after glitches 4, 5, and 6 exhibits a similar slope, which is approximately $\ddot{\nu}\sim 2.0\times 10^{-22}$ s-3.
We apply the vortex creep model to fit the three glitches, using Eqs. (7) and (14), and the $\chi_{\rm r}^{2}$ values are 1.09 (19 d.o.f.), 0.86 (18 d.o.f.), and 0.64 (12 d.o.f.) in order. The fit result is shown in Fig. 3 panel (d):
  • G4: The exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=86(9)$ d is slightly different than $\tau_{\rm d}=72(4)$ d found from fit with Eq. (2) (see Table 3) again due to the initial exponential dependency of Eq. (14). The vortex creep model prediction for the time to the next glitch from G4 is $t_{\rm th}=1680(16)$ d and in qualitative agreement with the observed inter-glitch time $t_{\rm obs}=1648(4)$ d. The total fractional crustal superfluid moment of inertia involved in G4 after corrected for the crustal entrainment effect is $I_{\rm cs}/I=2.30(2)\%$. The recoupling time-scale for the crustal superfluid associated with G4 is $\tau_{\rm nl}=3.4$ d.
@@ 4.4 PSR J1718--3825
PSR J1718--3825 is a young pulsar, discovered by Manchester et al. (2001) in the radio band. It has a characteristic age of 89.5 kyr, a spin period of 74.7 ms, and a relatively high spin-down luminosity of $\dot{E}\sim 1.3\times 10^{36}$ erg s-1. In addition to an associated X-ray nebula (Hinton et al., 2007), it is also a source of gamma-ray emission (HESS J1718) within the supernova remnant RX J1713.7--3946 (Aharonian et al., 2007).
We analysed more than 14 yr of timing data for PSR J1718--3825 in the radio and gamma-ray bands. A total of 3 glitches were identified, of which glitch 1 and 2 have been reported in the radio band (Yu et al., 2013; Lower et al., 2021), and glitch 3 is a new small glitch. By examining the timing residuals, it was deduced that glitch 3 occurred at MJD $\sim$ 59121(8). Using the fitting method, we determine the fractional glitch size to be $\Delta\nu/\nu\sim 1.9(2)\times 10^{-9}$ and the first frequency derivative to be $\Delta\dot{\nu}/\dot{\nu}\sim-0.12(9)\times 10^{-3}$. We also analyzed the timing noise using the pre- and post-glitch timing data of glitch 2. It was found that the spectral index remained consistently at 6 before and after the glitch.
In Fig. 5, the three detected glitches in PSR J1718--3825 are presented. It is clear that $\dot{\nu}$ only experienced a long-term linear recovery after the glitches, and there may be stable $\ddot{\nu}$. Therefore, based on the glitch epochs, the $\dot{\nu}$ data were divided into four intervals. and linear fits to $\dot{\nu}$ data were applied to each interval to determine $\ddot{\nu}$. The $\chi_{\rm r}^{2}$ values of these linear fits were 1.54, 0.66, and 0.74, respectively, with corresponding d.o.f. of 2, 30, and 12. The fits revealed that, under the influence of glitch, $\ddot{\nu}$ has undergone obvious changes, with values of $2.9(1)\times 10^{-23}$ s-3, $2.04(2)\times 10^{-23}$ s-3, and $1.65(9)\times 10^{-23}$ s-3, respectively. These fitted values are represented by the blue dashed lines in panel (c) of Fig. 5. A comparison with predictions from the superfluid vortex bending model is also provided for the first (G1) and second (G2) glitch events. Detailed discussions are included below in Sec. 5.
@@ 4.5 PSR J1509--5850
PSR J1509--5850 is a middle-aged pulsar with a characteristic age of 154 kyr (Kramer et al., 2003). We have analysed nearly 14 yr of timing data in both the radio and gamma-ray bands and found no glitches within the observed data range. Our study of the spin-down rate evolution suggests that PSR J1509--5850 exhibits a roughly stable $\ddot{\nu}$, as shown in Fig. 6. Using the least squares method for linear fitting of $\dot{\nu}$ data (yielding an $\chi_{\rm r}^{2}$ value of 0.80 with 19 d.o.f.), we determined $\ddot{\nu}$ to be $1.33(3)\times 10^{-24}$ s-3, with the fitted result indicated by the blue dotted line in Fig. 6. Additionally, the figure presents a fit based on the superfluid vortex bending model, which describes well the $\dot{\nu}$’s oscillation (see Sec. 5).
@@ 5.1 Oscillations in the spin-down rate
We apply the vortex bending model to the timing solution of PSR J1509--5850 between MJD 54303–59510, obtaining a fit result with a $\chi_{\rm r}^{2}$ value of 0.87 (15 d.o.f.), see Fig. 6. This result represents a significant improvement compared to the linear fit provided earlier in Sec. 4.5. The model free parameters corresponding to the vortex bending model’s application, along with the oscillation time $T_{\mathrm{osc}}$ and length scale $\ell$ are given in Table 6. If we attribute the large frequency second time derivative of $\ddot{\nu}\cong 1.37\times 10^{-24}$  s-3 to an unresolved Vela-like glitch occurred before the time interval we considered, then Eq. (24) with the typical value $\Delta\dot{\nu}_{\rm p}/\dot{\nu}=5\times 10^{-3}$ gives the estimate of $t_{\rm ig}=158$ yr for the time between two successive giant glitches hypothesised to be experienced by this pulsar. Thus, we predict that PSR J1509--5850 will glitch again within the same timescale. Future observations of this pulsar are required to test our assertion.
We also analyse the timing data of PSR J1718--3825 between MJD 54303–59491. Since there is not enough data for the third glitch (G3), we made an application of the vortex bending model to its first (G1) and second (G2) glitch events. See Fig. 5, panel (c). The fit parameters corresponding to the vortex bending model are given in Table 6. The $\chi_{\rm r}^{2}$ values for the fitting results of G1 and G2 are 0.76 (26 d.o.f.) and 0.17 (8 d.o.f.), respectively. We see that, while the vortex bending model slightly improves the fit to the data of G1 compared to the linear one (provided above in Sec. 4.4), it provides an excellent fit to the G2 data reducing the residuals significantly over the linear one. In particular, the damping timescale of the oscillations in the spin-down rate is well constrained to be $\tau=218$ d. When assessed with Eq. (8), this would provide important aspects for the microphysical parameters related to nuclear pinning mechanism as well as temperature estimate for the neutron star crust. Also, the length-scale $\ell$ found from the fits would provide a new method to put constraints on the crustal thickness and in turn on the EOS of the neutron stars. It should be also important to address these constraints in light of the sinusoidal-like oscillations observed in the spin-down rates of various pulsars.
@@ 5.2 Inter-glitch Braking Index
If only the external torque is considered, based on the use of torque laws of the form $-K\Omega^{n}$, where $K$ is a constant and the braking index is assumed to be 3 as predicted by the magnetic dipole model (Manchester & Taylor, 1977). Following the vortex creep model (Alpar et al., 1984; Gügercinoǧlu & Alpar, 2017), glitches may occur as a result of collective unpinning of a very large number of vortex lines in the crustal superfluid and inter-glitch spin evolution of a pulsar is dominated by the response of the star to the variations in the superfluid internal torques’ coupling due to such vortex discharges. In terms of the model parameters theoretical expectation of the inter-glitch braking index can be recast as (Alpar & Baykal, 2006):
@@
where the glitch magnitudes in the rotation and the spin-down rates are expressed in units of $10^{-6}$ and $10^{-3}$, respectively. Also, we have used the shorthand notation $\beta=I_{\rm B}/I_{\rm A}$. As a fiducial value, in our calculations below we take $\beta=4.11$, which was obtained for PSR J1048–5832 (Liu et al., 2024). In comparison, the $constant$ braking index ($n$) before and after the glitch were estimated according to the fitted $\ddot{\nu}$ and the parameters ${\nu}$, $\dot{\nu}$ in Table 2. We observed stable values of $\ddot{\nu}$ for PSRs J1028--5819, J1509--5850 and J1718--3825 and fitted the slopes of $\ddot{\nu}$ for each. In the case of PSR J1028--5819, the braking indices before and after the glitch are 21.7(4) and 63(1), respectively. For glitch 2 of PSR J1718--3825, the braking indices before and after the glitch are 49.2(5) and 40(2), respectively. Although we did not identify a glitch event for PSR J1509--5850, its $\dot{\nu}$ has steadily evolved over 14 yr, resulting in a braking index of $n=11.1(3)$. These non-canonical braking indices may be indicative of internal torque at work, likely associated with the dynamics of nucleon superfluidity in this context. Furthermore, PSR J1028--5819 exhibits a significant change in the braking index before and after the glitch, while maintaining a consistent spectral index ($\alpha$) of 4 in both periods (see Sec. 4.2). This suggests that the change in the braking index is unlikely to be related to red noise.
In addition, the most prominent effect of vortex unpinning on the spin parameters is $\ddot{\nu}$ and given by
@@
$t_{\rm ig}=\frac{\Delta\dot{\nu}_{\rm p}}{\ddot{\nu}_{\rm ig}}=2\times 10^{-3}\left(\beta+\frac{1}{2}\right)\tau_{\rm c}\left(\frac{\Delta\dot{\nu}_{\rm p}}{\dot{\nu}}\right)^{-1}_{-3}\left(\frac{\Delta\nu_{\rm p}}{\nu}\right)_{-6}\ .$ (24)
@@
Comparison of the prediction of the vortex creep model [Eq. (23)] with observed inter-glitch $\ddot{\nu}$ for our sample of glitches is given in Table 7. An inspection to the Table 7 reveals that the agreement between the vortex creep model predictions and observations are qualitatively good except for G8 of PSR J1420--6048. The reason for the discrepancy may be a persistent step increase that not resolved which leads to a somewhat larger estimate in $\Delta\dot{\nu}_{\rm p}$. Future observations will help to test this hypothesis.